# MEEG and AT-DGNN: Advancing EEG Emotion Recognition with Music and Graph Learning

### v1

@@ V RESULTS AND DISCUSSION
In this section, we compare the average accuracy and F1 score of AT-DGNN on the MEEG dataset with CNN, TCN, and GNN-based SOTA methods in the BCI domain. The CNN-based methods include: EEGNet [9], TSception [29], DeepConvNet and ShallowConvNet [10]. The TCN-based methods include: EEGTCNet [11], TCNet-Fusion [12] and ATCNet [30]. The GNN-based methods include: DGCNN [18] and LGGNet [16]. Due to the smaller size of the CNN-based models, the learning rate and the number of training epochs are reduced to avoid overfitting. For the other models, the parameters recommended by the authors are used. Additionally, ablation studies are conducted to reveal the contribution of each component within the AT-DGNN architecture.
@@ V-A Emotion Recognition
Table I lists the results of our proposed method compared to other SOTA methods on the MEEG dataset. The AT-DGNN series significantly outperforms the LGGNet series across both arousal and valence dimensions. Specifically, the AT-DGNN-Gen model achieve the highest F1 score in the valence dimension at 85.89%, which represents an improvement of 3.42% over the highest score of the LGGNet series (82.47%). In the arousal dimension, AT-DGNN-Fro leads with an accuracy of 84.88%, outperforming the best LGGNet result by 2.54%. The AT-DGNN models also demonstrate greater consistency with the smallest standard deviations, indicating robustness in the challenging and class-imbalanced MEEG dataset.
Compared to traditional CNN-based methods and other machine learning techniques, the AT-DGNN series show substantial enhancements. Improvements in F1 scores are up to 17.61% higher, and accuracy gains are as much as 9.66% better, highlighting the effectiveness of the DGNN approach. These improvements underscore the capability of AT-DGNN models to better capture complex patterns and dependencies in graph-structured data, thus providing more accurate and reliable performance across varied testing conditions.
The above methods all achieve relatively high accuracy rates on the MEEG dataset, with CNN-based models reaching around 75% accuracy, or even higher. In contrast, GNN-based models have accuracy rates above 80%. According to a study by Huang et al. [31] in 2023, the same models achieve about 60% accuracy on the DEEP dataset, which is significantly lower than on the MEEG dataset. This also demonstrates that music, compared to video, is more effective in evoking emotions in subjects.
@@ V-B Ablation Study
TABLE II: Ablation studies on Arousal of MEEG using AT-DGNN-Gen.
SW MHA D ACC(%) Changes(%) F1(%) Changes(%)
81.62 -1.01 82.16 -1.81
78.41 -7.16 79.99 -3.98
80.52 -2.11 81.22 -2.75
82.63 - 83.97 -
TABLE III: Ablation studies on Arousal of MEEG using AT-DGNN-Gen.
SW MHA D ACC(%) Changes(%) F1(%) Changes(%)
82.20 -3.17 81.46 -4.43
78.21 -7.68 78.93 -6.96
81.25 -4.12 81.10 -4.27
85.37 - 85.89 -
@@
SW: Sliding window segmentation.
MHA: Multi-head attention mechanism.
D: Dynamic graph neural network.
Changes: Compared with the baseline AT-DGNN-Gen model.
We conduct ablation studies to evaluate the contributions of different components of the AT-DGNN-Gen model on the arousal classification task using MEEG data. Table II and III presents the performance of the model with various combinations of its key components: sliding window segmentation (SW), multi-head attention mechanism (MHA), and dynamic graph neural network (D). The baseline model incorporates all three components.
Removing any single component results in a notable performance drop. For the A label, excluding SW, MHA, or D reduces the ACC by 1.01%, 7.16%, and 2.11%, respectively, and the F1 score by 1.81%, 3.98%, and 2.75%, respectively. For the V label, excluding SW, MHA, or D reduces the ACC by 3.17%, 7.68%, and 4.12%, respectively, and the F1 score by 4.43%, 6.96%, and 4.27%, respectively. These findings highlight the importance of each component in enhancing the model’s ability to capture and utilize complex data relationships, validating the architectural design of AT-DGNN.
@@ VI Conclusion
This study introduces the MEEG dataset and the AT-DGNN framework as significant advancements in EEG-based emotion recognition. The MEEG dataset, which uses music to induce emotional states, provides a valuable resource for studying brain responses to emotional stimuli. The AT-DGNN framework, integrating attention mechanisms with DGNN, excels in capturing complex temporal brain dynamics, significantly improving emotion recognition accuracy compared to other SOTA methods. These innovations advance brain-computer interface technologies and open new avenues for exploring the relationship between music, emotion, and brain activity. The superior performance of AT-DGNN, validated against traditional models, underscores its potential for enhancing emotion analysis.

### v2

@@ V RESULTS AND DISCUSSION
In this section, we compare the average ACC and F1 score of AT-DGNN on the MEEG dataset with CNN, TCN, and GNN-based SOTA methods in the BCI domain. The CNN-based methods include: EEGNet [9], TSception [24], DeepConvNet and ShallowConvNet [10]. The TCN-based methods include: EEG-TCNet [11], TCNet-Fusion [12] and ATCNet [31]. The GNN-based methods include: DGCNN [18] and LGGNet [16]. Due to the smaller size of the CNN-based models, the learning rate and the number of training epochs are reduced to avoid overfitting. For the other models, the parameters recommended by the authors are used. Additionally, ablation studies are conducted to reveal the contribution of each component within the AT-DGNN architecture.
@@ V-A Emotion Recognition
As shown in Table I, the AT-DGNN series, particularly the AT-DGNN-Gen model, significantly outperforms the previous SOTA LGGNet series in both arousal and valence dimensions, with a notable increase in F1 scores and accuracy. The AT-DGNN-Gen model not only achieves the highest accuracy of 83.74% in the arousal dimension, surpassing the LGGNet-Gen model by 1.59%, but also records the highest accuracy of 86.01% in the valence dimension, a substantial improvement of 3.48% over the LGGNet series. This improvement underscores the efficacy of our approach in capturing complex patterns within the MEEG dataset, which is known for its challenging and class-imbalanced nature.
The AT-DGNN-Gen model exemplifies the superior performance of the AT-DGNN series, achieving robust results with an average accuracy of 84.88% and an F1 score of 85.12% across emotion dimensions. This underscores the effectiveness of our graph-based approach in handling complex emotion recognition tasks. Furthermore, consistent with the findings of [16], the graph configuration $G_{g}$ demonstrated superior performance in both AT-DGNN and LGGNet models compared to other graph definitions, indicating its optimal structure for capturing relevant features in emotion analysis.
Furthermore, these results demonstrate a marked improvement over traditional CNN and TCN-based methods, with average gains in F1 scores and accuracy surpassing previous models by up to 7.16% and 5.78%, respectively. This emphasizes the advanced capability of the DGNN approach in graph-structured data analysis, providing more accurate and reliable emotion recognition.
All methods achieve high accuracy on the MEEG dataset, with CNN and TCN models reaching 75% or higher, and GNN models exceeding 80%. On the DEAP dataset, AT-DGNN averages 60.56% accuracy across four dimensions. According to Huang et al. [32], other models, including LGGNet, also achieve around 60% accuracy on DEAP, with some slightly outperforming AT-DGNN and others not. All models perform significantly better on the MEEG dataset, suggesting that music is more effective than video in evoking emotions.
@@ V-B Ablation Study
TABLE II: Ablation study on Emotion of MEEG using AT-DGNN-Gen.
S M D ACC(%) Changes(%) F1(%) Changes(%)
81.91 -2.97 81.81 -3.31
3 78.31 -6.57 79.46 -5.66
3 80.89 -3.99 81.16 -3.96
1 81.91 -2.97 81.81 -3.31
2 82.94 -1.94 82.52 -2.60
4 83.17 -1.71 83.48 -1.64
3 84.88 - 85.12 -
  • S: Sliding window segmentation.
  • M: Multi-head attention mechanism.
  • D: Dynamic graph neural network.
@@
  • ✗: Exclude the component.
  • 1, 2, 3, 4: The number of stacked DGNN layers.
  • Changes: Compared with the baseline AT-DGNN-Gen.
As illustrated in Table II, an ablation study is conducted to systematically assess the individual contributions of critical components within the AT-DGNN model: sliding window segmentation (S), multi-head attention mechanism (M), and the dynamic graph neural network (D). The results underscore the pivotal roles of these components, indicating that their omission results in declines in accuracy (ACC) of 2.97%, 6.57%, and 3.99% and in F1 scores of 3.31%, 5.66%, and 3.96%, respectively, thereby demonstrating their functionality and synergy.
Furthermore, the experiment demonstrated that the number of layers in the DGNN significantly affects the model’s performance. Modifying the number of DGNN layers to $1$, $2$, or $4$ led to reductions in accuracy of 2.97%, 1.94%, and 1.71%, and in F1 scores of 3.31%, 2.60%, and 1.64%, respectively. A configuration with a single layer implies non-utilization of the stacked DGNN, highlighting the importance of network depth tuning in optimizing the model’s capacity to process and analyze complex relationships in EEG data.
The optimal configuration, employing $3$ layers of the DGNN, is found to be the most effective, as evidenced by achieving the highest scores in both ACC and F1 metrics. These findings substantiate the design choices of the AT-DGNN model and shed light on the interaction among its components, significantly enhancing computational neuroscience analysis.
@@ VI Conclusion
This study introduces the MEEG dataset and the AT-DGNN framework, marking substantial progress in the field of EEG-based emotion recognition. The MEEG dataset uniquely employs music to induce emotional states, providing an invaluable resource for probing brain responses to emotional stimuli. The AT-DGNN, which integrates attention mechanisms with DGNN, adeptly captures the intricate temporal dynamics of brain activity, thereby markedly enhancing emotion recognition accuracy over existing SOTA methods. This advancement not only propels BCI technology forward but also facilitates novel investigations into the interplay among music, emotion, and cerebral activity. The exemplary performance of AT-DGNN, validated through rigorous comparisons with traditional models, solidifies its utility in advancing neuropsychological understanding.